ECAL: Problematic channels / actions log

Section/X/Y Time Description Comment
31 operational cells, which were masked to fight noisy lines 04.09.2008, ~11:30 Unmasked back i.e. nominal PM HV values in DB are recovered
LED group A103 03.09.08, ~16:00 RJ-45 connector is plugged to spare socket of the same LEDTSB unit. Absent signal from A103.3. Database is updated 04.09.2008, ~11:00 am
LED group A105 03.09.08, ~16:00 RJ-45 connector is plugged to spare socket of the LEDTSB unit in crate 15 / slot 1. Absent signal from A105.1. Database is updated 04.09.2008, ~11:00 am
Outer section periphery 03.09.08, ~16:00 ~80 patches (total amount) with 45mkF filtering capacitances are installed in the output of control boxes. C-side: voltage control / distribution boxes 1, 2, 3, 4, 8, 9 (partially) except for 2 lines (O 11/32..57 – bad patch, removed after first MV switching ON; O 14/32..57 – patch is missing? to be checked). A-side: voltage control / distribution boxes 32, 31, 30, 29, 25
LED group C117 29.08.08, ~16:00 RJ-45 connector is plugged to spare socket of the same LEDTSB unit. Mapping change: C117.1: 13/00/08 -> 13/00/56; C117.2: 13/00/11 ->13/00/59; C117.3: 13/00/09 -> 13/00/57; C117.4: 13/00/10 -> 13/00/58 Absent signals from C117.2/4. Database is updated 30.08.2008, ~07:00 am
Power Lines O/63/06..31 and O/60/06..31 29.08.08, ~16:00 Patches with 45mkF filtering capacitances are installed in the output of control box Test batch to study suppression of HF modulation of MV level
LED group C116 20.08.08, around noon RJ-45 connector is replaced Absent signal from C116.4 / unstable signal from C116.1
All ECAL cells 17.08.08 10:30 Nominal settings for ECAL HV were changed in Calo DB to respect the spread in PM quantum efficiencies Clipping coeff: 1/3; New formula to relate PM control voltage and high voltage in the C-W output HV (V) = 38.519 + 432.695 * PmControl (V)
LEDTSB 13/0 14.08.08 Replaced with another one Problem with delay chips configuration
LED group C213 24.07.08 RJ-45 is reconnected to the spare socket on the same LEDTSB unit. Mapping change: C213.1: 13/01/49 -> 13/01/57; C213.2: 13/01/51 -> 13/01/59; C213.3: 13/01/48 -> 13/01/56; C213.4: 13/01/50 -> 13/01/58 No trigger pulse for LED C213.2. Database is updated at the same time.
A-side, voltage distribution/control boxes 07-...06.08 Modification: 100 Ohm in MV chain + ordinary fuses in LV chains to shunt the resettable ones  
M/54/12...M/54/51 02.07.08 evening - 03.07.08 14:45 Short circuit in power line 27/7 (-6V LV chains). Provoked by Xavier's tests of correct MV shut-down before switching off LV Line is re-plugged to spare 25/0. Mapping in DB is updated accordingly at the same time
C-side, voltage distribution/control boxes 15-21.06.08 Modification: 100 Ohm in MV chain + ordinary fuses in LV chains to shunt the resettable ones  
O/12/32..O/12/57 02.06.08 afternoon Connector on the power line (4/0) replaced 3 June ~10:00. Bad connection quality
O/14/06..O/14/31 Connector on the power line replaced 3 June ~10:00. Line is connected to 8/0 (! previous one is 8/6) Bad connection quality. Mapping in DB is updated accordingly 3 June 2008, around 14:00
O/14/06..O/14/31 02.06.08 ~12:00 Reconnected from 4/5 to the spare 8/6 Showed short circuit in +6V (all power supplies were on). After immediate turning OFF of Agiltent/MARATON fuse successfully recovered in one hour, but showed resistance 3.6 Ohm. Mapping in DB is updated accordingly 2 June 2008, around 14:00
O/15/32..O/15/57 Reconnected from 4/6 to the spare 8/7 During the check of the fuses on the box #4 showed resistance 3.6 Ohm. Mapping in DB is updated accordingly 2 June 2008, around 14:00
O/14/32..O/14/57 Box 4 / connector 5 - to be watched During the check of the fuses on the box #4 showed resistance 3.2 Ohm
ECAL A-side 02.06.08 ~10:00 Multiple short circuits in LV +6V Certainly provoked by somebody's activity, since OV protection relay states had been changed also
Voltage distribution / control box 31 31.05.08 13:00 Replaced with the spare one Positive LV seems to be unstable in all output connectors. Old serial number 46, new serial number 43. DB is updated accordingly 31 May 2008, 13:00
M/52/12...M/52/51 31.05.08 13:00 Fuse in 26/1 had been replaced, therefore line is plugged back from 25/5 to 26/1. Mapping in DB is updated accordingly 31 May 2008, evening
O/38/06...O/38/57 29.05.08 ?:?? - 31.05.08 13:00 Short circuit in power line 24/4 (+6V LV chains) Damaged fuse. Line is re-plugged to spare 24/7. Mapping in DB is updated accordingly 31 May 2008, evening
M/09/12...M/09/51 26.05.08 ?:?? - 28.05.08 ??:?? Short circuit in power line 6/7 (+6V LV chains) Damaged fuse, replaced
O/20/06...O/20/57 Short circuit in power line 8/4 (+6V LV chains)
I/45/14...I/45/49 26.05.08 ?:?? - 27.05.08 ??:?? Short circuit in power line 21/0 (+6V LV chains)
M/53/12...M/53/51 Short circuit in power line 26/0 (+6V LV chains)
M/55/12...M/55/51 Short circuit in power line 27/6 (+6V LV chains)
M/54/12...M/54/51 Short circuit in power line 27/7 (+6V LV chains)
M/63/02...M/63/51 Short circuit in power line 28/0 (+6V LV chains)
M/62/12...M/62/51 Short circuit in power line 28/1 (+6V LV chains)
O/43/06...O/43/57 Short circuit in power line 25/4 (+6V LV chains)
26.05.08 Power Cut. Agilent's stayed ON while MARATONS were OFF (~1hour)
MARATON C-side channel 4 22.05.08 morning - 22.05.08 17:00 Was in error state. Last current reading 23A(!) Simple switching OFF/ON helped. Consumptions checked, OK
O/34/6...O/34/57 ? 21.05.08 16:30 - 22.05.08 20:00 Abnormally high current consumption from AGILENT provokes trip of OV protection relays Malfunctioning C-W base (cell O 34/19), replaced. PM id and HV setting are updated in DB 23 May 2008, 15:00
M/52/12...M/52/51 ? 21.05.08 16:30 - 21.05.08 21:00 Short circuit in power line 26/1 (+6V LV chains) Damaged fuse. Line is re-plugged to spare 25/5. Mapping in DB is updated accordingly 23 May 2008, 15:10
21.05.08, ~16:00 Massive OV protection faults over all ECAL (both sides) for reason unknown
16.05.08, afternoon Local Power Cut, A-side
09.05.08 Power Cut (?rumors)
All Ecal channels 08.05.08 11:00 Nominal settings for ECAL HV were changed in Calo DB to follow the rule [ 10 + 7. x sin(theta) ] GeV of transverse energy for an upper limit of ECAL dynamic range The earlier rule was 10GeV of transverse energy
30.04.08 ~11:20 Power Cut
O/60/32...O/60/57 29.04.08 8:50 - 01.05.08 14:00 Short circuit in power line 32/6 (+6V LV chains) Fixed (malfunctioning C-W base replaced)
17.04.08, morining Power Cut (short)
I/19/14...I/19/49 09.04.08 ?:?? - 11.04.08 ??:?? Short circuit in power line 12/1 (+6V LV chains) Damaged fuse, replaced
M/00/12...M/00/51 Short circuit in power line 5/0 (+6V LV chains)
M/01/12...M/01/51 Short circuit in power line 5/1 (+6V LV chains)
M/08/12...M/08/51 Short circuit in power line 6/6 (+6V LV chains)
M/10/12...M/10/51 Short circuit in power line 7/0 (+6V LV chains)
08.04.08 Power Cut (?rumors)

-- IrinaMachikhiliyan - 20 Aug 2008

Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2008-09-04 - IrinaMachikhiliyan
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback